Midlands Inc. had a bad year in 2016. For the first time in its history, it operated at a loss. The company’s income statement showed the following results from selling 77,000 units of product: net sales $2,310,000; total costs and expenses $1,944,000; and net loss $-366,000. Costs and expenses consisted of the following.

Total

Variable

Fixed

Cost of goods sold $1,275,000 $774,000 $501,000
Selling expenses 520,000 94,000 426,000
Administrative expenses 149,000 56,000 93,000
$1,944,000 $924,000 $1,020,000


Management is considering the following independent alternatives for 2017.

1. Increase unit selling price 25% with no change in costs and expenses.
2. Change the compensation of salespersons from fixed annual salaries totaling $197,000 to total salaries of $40,000 plus a 5% commission on net sales.
3. Purchase new high-tech factory machinery that will change the proportion between variable and fixed cost of goods sold to 50:50.


(a) Compute the break-even point in dollars for 2017.

AND


(b) Compute the break-even point in dollars under each of the alternative courses of action.

1. Increasing selling price.

2. Change compensation.

3. Purchase machinery.

AND

Which course of action do you recommend?
Answer #1

(a)

Sales Revenue $         23,10,000
Variable Expenses $           9,24,000
Contribution Margin $         13,86,000
Fixed Expenses $         10,20,000
Net Operating Income $           3,66,000


Contribution Margin ratio = Contribution Margin / Sales x 100
= $1386000/2310000 = 60%

Break Even Point Dollars = Fixed Expenses / Contribution Margin ratio
= $1020000 / 60% = $1700000

(b)
1.

Sales Revenue $         28,87,500 =2310000*1.25
Variable Expenses $           9,24,000
Contribution Margin $         19,63,500
Fixed Expenses $         10,20,000
Net Operating Income $           9,43,500

Contribution Margin ratio = Contribution Margin / Sales x 100
= $1963500/2887500 = 68%

Break Even Point Dollars = Fixed Expenses / Contribution Margin ratio
= $1020000 / 68% = $1500000

2.

Sales Revenue $         23,10,000
Variable Expenses $         10,39,500 =924000+2310000*5%
Contribution Margin $         12,70,500
Fixed Expenses $           8,63,000 =1020000-197000+40000
Net Operating Income $           4,07,500

Contribution Margin ratio = Contribution Margin / Sales x 100
= $1270500/2310000 = 55%

Break Even Point Dollars = Fixed Expenses / Contribution Margin ratio
= $863000 / 55% = $1569091

3.

Sales Revenue $         23,10,000
Variable Expenses $           7,87,500 =924000-774000+1275000*50%
Contribution Margin $         15,22,500
Fixed Expenses $         11,56,500 =1020000-501000+1275000*50%
Net Operating Income $           3,66,000

Contribution Margin ratio = Contribution Margin / Sales x 100
= $1522500/2310000 = 65.91%

Break Even Point Dollars = Fixed Expenses / Contribution Margin ratio
= $1156500 / 65.91% = $1754690

Increasing unit selling price is recommended since it gives lowest break even point